Suppose the demand equation for cheese is


QD = 50 – 0.55P


Determine the price elasticity of demand when the price of cheese is

$6.00


$8.00


$10.00


$12.00


Identify whether demand is elastic, inelastic, or unit elastic at each price level.


1
Expert's answer
2021-09-12T19:33:11-0400

(a) Let's find the quantity of the product if price is $6.00:


"Q_d=50-0.55\cdot6=46.7."

Let's also write the inverse demand function:


"P=90.9-1.81Q_d."

We can find the price elasticity of demand when the price of cheese is $6.00 as follows:


"E_d=\dfrac{-1}{slope}\cdot\dfrac{P}{Q_d},""E_d=\dfrac{-1}{-1.81}\cdot\dfrac{6}{46.7}=0.07."

Since "E_d<1", the price elasticity of demand is inelastic at this price level.

(b) Let's find the quantity of the product if price is $8.00:


"Q_d=50-0.55\cdot8=45.6."

Then, we can find the price elasticity of demand when the price of cheese is $8.00 as follows:


"E_d=\dfrac{-1}{-1.81}\cdot\dfrac{8}{45.6}=0.09."

Since "E_d<1", the price elasticity of demand is inelastic at this price level.

(c) Let's find the quantity of the product if price is $10.00:


"Q_d=50-0.55\cdot10=44.5."

Then, we can find the price elasticity of demand when the price of cheese is $10.00 as follows:


"E_d=\dfrac{-1}{-1.81}\cdot\dfrac{10}{44.5}=0.12."

Since "E_d<1", the price elasticity of demand is inelastic at this price level.

(d) Let's find the quantity of the product if price is $12.00:


"Q_d=50-0.55\cdot12=43.4."

Then, we can find the price elasticity of demand when the price of cheese is $12.00 as follows:


"E_d=\dfrac{-1}{-1.81}\cdot\dfrac{12}{43.4}=0.15."

Since "E_d<1", the price elasticity of demand is inelastic at this price level.